A buffer is prepared using HLac and NaLac. 0.300dm3 of the 0.500mol/dm3 HLac solution is mixed with 0.300 dm3 of the 0.300mol/dm3 NaLac solution to prepare the buffer. Ka for HLac is 1.4×10^-4
1
Expert's answer
2020-09-29T06:48:50-0400
"pH = pK_a - log{c(HLac) \over c(NaLac)}"

"pH = 3.85 - log{0.300*0.500 \over 0.300*0.300} = 3.85-0.22 = 3.63"